Given two regu

lar polygons of the same number of sides. If the area of the first is

3 times the area of the second, find the ratio of (a) their perimeters, and (b) their

apothems.



1
Expert's answer
2021-01-28T05:21:13-0500

If two regular polygons of the same number of sides are given then such polygons are

similar


"k=\dfrac{side_1}{sie_2}=\dfrac{perimeter_1}{perimeter_2}=\dfrac{apothem_1}{apothem_2}"

"=\sqrt{\dfrac{Area_1}{Area_2}}"

Given "\dfrac{Area_1}{Area_2}=3"

(a)

"\dfrac{perimeter_1}{perimeter_2}=\sqrt{3}"

(b)


"\dfrac{apothem_1}{apothem_2}=\sqrt{3}"